OpenStudy (anonymous):

1. Find the minimum, maximum, range, mean, and median for the data set. 12, 13, 13, 17, 18, 19, 27 A. minimum: 12 maximum: 27 range: 15 mean: 17 median: 13 B. minimum: 27 maximum: 12 range: 15 mean: 17 median: 17 C. minimum: 12 maximum: 27 range: 15 mean: 17 median: 17 D. minimum: 12 maximum: 27 range: 17 mean: 15 median: 17 help me

OpenStudy (anonymous):

min =12 max=27 Range=max.-min.=27-12=15 Median=arranging in ascending order and Med= average of n/2and(n/2)+1,if n is even. =(n+1)/2,if nis odd. Heren=7,(no. of terms) so median=(7+1)/2=4th term=17 \[average=\frac{ 12+13+13+17+18+19+27 }{7}=?\] calculate.